# Prototyping a High Purity Germanium cryogenic veto system for a bolometric detection experiment

### v1

@@ Abstract
The use of High Purity Germanium detectors operated in ionization mode at cryogenic temperatures is investigated as an external background mitigation solution for bolometers used in rare-event search experiments. A simple experimental setup with very partial coverage, running a 52-g bolometer sandwiched in-between two 2-cm thick High Purity Germanium cylindrical detectors in a dry cryostat, shows promising rejection to environmental gammas and atmospheric muons backgrounds. The acquired data are used together with a Monte Carlo simulation of the setup to extract the main contributions to the external backgrounds expected in an above ground experiment, such as e.g. current and future experimental efforts targeting the detection of coherent elastic neutrino-nucleus scattering at reactor facilities. Based on all these results, a coverage similar veto system achieving a (10 keV) energy threshold is expected to achieve a 70 % and a 97 % rejection power for gamma-like and muon-like events, respectively.
@@ 1 Introduction
Over the past decades, the steady progresses achieved in the development of cryogenic detectors for particle and nuclear physics widespread their use in rare-event search experiments. The high energy resolution, the large sensitivity range and the possibility to implement multiple read-out strategies using heat, scintillation, and ionization signals make them extremely attractive and suitable devices for e.g. dark matter searches Cresst ; EdelweissTES ; SuperCDMS or searches for neutrino-less double beta decay Cupid-Mo . Although these detectors can already be very effective in identifying and discriminating unwanted sources of background, the quest for ever-higher sensitivities in the future generation of bolometric detection experiments call for innovative background rejection strategies pursuing for the highest efficiencies. This feature is especially relevant for the shortcoming generation of experiments aiming at studying coherent elastic neutrino-nucleus scattering (CENS) at reactors using cryogenic detectors, such as e.g. Nucleus NUCLEUS2019 or Ricochet Ricochet . The detection of CENS-induced nuclear recoils down to sub-keV energies still requires effective background rejection techniques in these uncharted very low energy regimes, and in which the ionization or scintillation yield degrades, making an effective background mitigation strategy solely based on a dual-readout scheme very challenging to achieve. To this end, this article investigates for the first time the feasibility of using a compact cryogenic veto system in close vicinity to a target cryogenic detector for mitigating external particle backgrounds. The use of High Purity Germanium (HPGe) ionization detectors, which are known to have a rather low (keV) energy threshold and fast timing response, is here considered.
@@ 2 Experimental setup
The experimental setup is pictured and sketched on figure 1. It consists of two cylindrical 400 g HPGe crystals of 70 mm diameter 20 mm height, each placed above and below a 25 mm diameter 25 mm height 51.7 g cylindrical crystal BASKET ; LWO_Growth .
@@
The two HPGe crystals were operated in ionization mode using Aluminum (Al) electrodes evaporated on their respective top and bottom surfaces to collect the electron/hole charge carriers. These 200-nm thick electrodes were deposited by electron beam evaporation after an Argon (Ar) cleaning and passivation step using a 30-nm thick amorphous Ge:H layer Evap . The bottom crystal (BOT Ge) is equipped with planar electrodes. The top crystal (TOP Ge), which was previously used in the Edelweiss experiment Edelweiss2010 , is equipped with ring-shaped electrodes (ID electrodes). These electrodes were specially designed to discriminate between surface and bulk events. Because this feature was unnecessary in the present work, the ring electrodes were serially connected together to mimic as closely as possible the electric field resulting from polarized planar electrodes such as those equipping the BOT Ge. The two HPGe crystals were housed within 3-mm thick copper boxes and secured using PTFE (Teflon) holders. The thermal link and the electrode signal transmission were ensured by bonding the crystals using a Kapton printed circuit board glued onto the copper box. The copper housing lids facing the LWO detector were mechanically designed to hold the LWO crystal using PTFE pieces. The HPGe detectors were polarized with a 5 V/cm electric field by biasing one electrode at +10 V and grounding the opposite one. An infra-red LED (EOLS-1650-199 from EPIGAP) was glued to the BOT Ge housing in order to regenerate the HPGe crystals regen_Ge . The two crystals were daily illuminated for ten minutes, after which one hour was necessary to stabilize the cryostat back to base temperature. A Am source was mounted underneath the BOT Ge detector in order to study its behavior at low energies close to the threshold. The copper housing was therefore drilled with a 2-mm diameter hole to allow the source to irradiate the crystal. The source was covered by a thin copper tape and a collimator to shield the detector against particles emitted in the Am decay.
@@
The LWO crystal was operated as a bolometer and read out with a Neutron-Transmutation-Doped (NTD) thermal sensor NTD . It was used as a target detector to test the identification and the rejection power of external particles using the HPGe ionization detectors operated in anti-coincidence mode. The NTD sensor was biased with a 1-kHz alternative constant current () in order to improve the signal-to-noise ratio NTD-Modulation . The full mechanical assembly was wrapped in a copper foil to shield against possible infra-red radiations. The experimental setup was operated at a temperature of 18 mK in the so-called "Actuator" cryostat at Laboratoire de Physique des Deux Infinis Irène Joliot-Curie (IJCLab, Orsay), which is a custom-made pulse-tube dilution refrigerator embedded in a single vacuum chamber (see e.g. thesis_beatrice for further details). No vibration decoupling systems were used.
@@
The detectors were connected with low noise stainless steel cables. The electronic operation and acquisition scheme is detailed on figure 2 for both the HPGe ionization detectors and for the LWO bolometer. A first signal amplification step is performed in a cold stage using Junction Field Effect Transistors (IFN860 bi-JFETs). The output voltage signals are then filtered and further amplified through voltage amplifiers in a second electronic stage operated at room temperature. Signal acquisition was done with a commercial NI6366 X Series Data Acquisition system from National Instruments. Output signals were continuously sampled and read out at a 100 kHz frequency.
@@
Data were collected in two experimental configurations: (i) without the use (10 hours of acquisition time) and (ii) with the use (40 hours of acquisition time) of an external lead (Pb) shielding (see right panel of figure 1) placed around the cryostat. The purpose of this shielding was to limit the pile-up of events in the slow LWO detector. It was made of three 5-cm thick brick walls and one partial wall of 10-cm thickness, hence partially covering the detector setup. During those two data acquisition runs, the detector setup has punctually been exposed to a source and a neutron source.
@@ 3 Experimental results
The operation of HPGe ionization detectors as an efficient veto system for a bolometric target detector imposes good energy and time reconstruction together with a high pulse selection efficiency down to the lowest possible threshold. This latter requirement is particularly relevant for veto detectors operated at surface, for which event rates can lead to significant pile-up probability. While a good timing resolution allows to keep accidental coincidences at acceptable rates, precise energy determination may prove to be important to better identify potential background sources present in the experimental setup environment. To meet these requirements, a fast and efficient pulse reconstruction algorithm was developed. It takes advantage of the large difference between the detector pulse rise and decay time components, then making it almost insensitive to the presence of pile-up events. In this situation, the information about the deposited energy is mostly contained in the rise time region of the pulse. Event triggering and energy determination can then be performed in a rather small time window, hence keeping under control possible pulse distortions from pile-up effects. The present event selection and reconstruction algorithm have been applied to both the HPGe veto detectors and the LWO bolometer data streams. It is described in the following subsections.
3.1 Characterization of the HPGe detectors
@@
Figure 3 illustrates typical signals recorded over a 1-s time window after combining the ionization signals from the TOP Ge detector anode and cathode. Signals are here expressed in volts, and are sampled at a 100 kHz frequency for their later processing and analysis. The amplified signals from the Ge detectors were measured to have typically an average rise time less than 30s and a decay time typically in the 30 to 50 ms range. Such fast rise times as compared to much slower decay times allow to perform a bin-to-bin voltage difference such that physical pulses can clearly be identified above a flat baseline centered around zero (see red plot on figure 3). Events are thus triggered by setting an appropriate threshold above the measured baseline fluctuations. The energy associated to each detected pulse is then reconstructed by summing the discrete derivative amplitudes over a time window covering the pulse leading edge region (i.e. non-zero derivative), and after applying energy calibration. Such a procedure is equivalent to measuring the full amplitude rise of the detector signal. In order to reduce baseline fluctuations and to improve energy resolution, event reconstruction was performed using an effective 50 kHz sampling rate by averaging amplitudes over time periods. In addition, the baseline level, measured before and after the triggered signal, was determined by averaging amplitudes over a longer range of .
Energy calibration was performed using a set of intense gamma-ray lines originating from the decay of primordial radionuclides naturally present in the surrounding environment: ( decay chain) at 510.7 keV, 583.2 keV and 2614.5 keV, annihilation at 511 keV, ( decay chain) at 609.3 keV, ( decay chain) at 911.2 keV and 969.0 keV, and at 1460.8 keV. An additional 59.5 keV gamma-ray line from the source was also used for the BOT Ge detector. Figure 4(a) shows the obtained calibration curves (blue and red dashed lines). Taking into account the gain of 1000 provided by the amplification stage, the BOT Ge detector featuring planar electrodes is found to have a nV/keV sensitivity while a nV/keV sensitivity is measured for the TOP Ge detector having the ring-shaped electrodes. Non-linearities in the energy scale below 3 MeV are found to be negligible. Relative energy resolutions measured at the abovementioned gamma-ray line energies are displayed on Figure 4(b). They are found to be better than 1 % above 400 keV and are very similar for both detectors. They are well described by the characteristic energy dependence as expected from Ge semiconductor detectors knoll :
@@
Electronic noise contribution for each of the HPGe detector was independently characterized by sampling the baseline fluctuations during each data run. The mean of the distribution of the baseline r.m.s. values were measured to be 1.1 keV and 1.8 keV for the BOT and TOP Ge detectors, respectively. These measurements come in very good agreement with the previously mentioned fit results (see figure 4 (b)). The larger value, in keV units, observed for the TOP Ge detector is mostly due to its smaller sensitivity. Some spurious periodic pulses were also identified and always reconstructed with energies below 15 keV. After investigation, those were attributed to cross-talk effects originating from the 1-kHz modulation current flowing through the LWO NTD sensor. Since the sampling and modulation clocks were synchronous, this cross-talk noise could be mitigated by removing very low energy pulses that were identified in-time with the leading edges of the NTD current modulation signal. This requirement introduced a small inefficiency of at most a few percents in the selection of events below 15 keV.
With the noise treatment described above, detection thresholds of 5 keV and 10 keV were applied to select and reconstruct events in the BOT and TOP Ge detectors, respectively. The corresponding counting rates, integrated over the full energy spectra up to 20 MeV, were measured to be about 15 and 8 counts/s, respectively, during a typical background run using the Pb shielding. The difference in rate between the two detectors can be mostly accounted for by the activity of the source. Figure 5 further illustrates this difference, where the energy spectra of the selected and reconstructed events in the TOP and BOT Ge detectors are compared over the full 20 MeV range and in the region below 600 keV. With no Pb shielding surrounding the cryostat, the rates were found to be significantly higher, ranging between 37-43 and 26-31 counts/s for the BOT and TOP Ge detectors, respectively.
@@
As mentioned above, good time resolution of the Ge veto detectors is necessary in order to limit accidental rejection of good candidates in the LWO bolometer. The arrival time of an event in the BOT and TOP Ge detectors was determined to better than 10 s, owing to the short rise time of pulses. Given the counting rates, true coincidences between the two Ge detectors could clearly be observed with a negligible level of accidentals. This is illustrated also in Figure 5, where the TOP and BOT Ge energy spectra are also shown for events detected in coincidence (orange and green curves). With the exception of the 511 keV annihilation line, all other X-ray and gamma-ray lines corresponding to full absorption peaks disappear. Moreover, a large fraction of events detected in coincidence is located in the high energy portion of the spectra. These events come from atmospheric muons crossing both detectors. The rate of coincidences bewteen the TOP and BOT Ge detectors for a typical background run using the Pb shielding is 0.8 counts/s.
3.2 LWO bolometer characterization
A similar analysis was carried out for the selection and energy reconstruction of events in the LWO bolometer. Since NTD pulses have here a much longer rise and decay times, typically 25-ms and up to 1 s, respectively, an effective sampling frequency of 400 Hz was applied by averaging amplitudes over 2.5 ms time periods in the initial streams of data. This choice results from a trade-off between energy and time resolution. In order to reject possible spurious events, pulses with less than 10-ms rise time were rejected. Figure 7 shows an example of pulses from the LWO bolometer, recorded over a 10-s time window.
The selection and energy determination of events were performed using the same technique based on the bin-to-bin amplitude subtraction. Figure 8 shows the energy spectrum measured with the LWO bolometer during a run performed with the Pb shielding. In addition to environmental gamma-ray lines, a peak at MeV due to the neutron capture reaction is visible. The peak intensity is here enhanced by the use of a neutron source. It is shifted by with respect to the 4.78 MeV total energy released in the reaction. This effect corresponds to the so-called thermal quenching, which is expected to be of the order of a few percents in Lithium-containing scintillating bolometers BASKET .
The energy calibration of the LWO bolometer follows the exact same principles as the procedure described in the previous section, and was obtained using the most prominent gamma-ray lines observed below 3 MeV (see Figure 4 (a)). Taking into account a gain of 8000 from the NTD signal warm amplification stage, the sensitivity of the bolometer was determined to be nV/keV. The measured energy resolution below 6 MeV is shown on figure 4 (b). It is well-described by the following relation Armengaud_2017 ; Augier_2022 :

### v2

@@ Abstract
The use of High Purity Germanium detectors operated in ionization mode at cryogenic temperatures is investigated as an external background mitigation solution for bolometers used in rare-event search experiments. A simple experimental setup, running a 52-g bolometer sandwiched in-between two 2-cm thick High Purity Germanium cylindrical detectors in a dry cryostat, shows promising rejection to environmental gammas and atmospheric muons backgrounds. The acquired data are used together with a Monte Carlo simulation of the setup to extract the main contributions to the external backgrounds expected in an above ground experiment, such as e.g. current and future experimental efforts targeting the detection of coherent elastic neutrino-nucleus scattering at reactor facilities. Based on all these results, a coverage similar veto system achieving a (10 keV) energy threshold is expected to achieve a 73 % and a 92 % rejection power for gamma-like and muon-like events, respectively.
@@ 1 Introduction
Over the past decades, the steady progresses achieved in the development of cryogenic detectors for particle and nuclear physics widespread their use in rare-event search experiments. The high energy resolution, the large sensitivity range and the possibility to implement multiple read-out strategies using heat, scintillation, and ionization signals make them extremely attractive and suitable devices for e.g. dark matter searches [1, 2, 3] or searches for neutrino-less double beta decay [4]. Although these detectors can already be very effective in identifying and discriminating unwanted sources of background, the quest for ever-higher sensitivities in the future generation of bolometric detection experiments call for innovative background rejection strategies pursuing for the highest efficiencies. This feature is especially relevant for the shortcoming generation of experiments aiming at studying coherent elastic neutrino-nucleus scattering (CENS) at reactors using cryogenic detectors, such as e.g. Nucleus [5] or Ricochet [6]. The detection of CENS-induced nuclear recoils down to sub-keV energies still requires effective background rejection techniques in these uncharted very low energy regimes, and in which the ionization or scintillation yield degrades, making an effective background mitigation strategy solely based on a dual-readout scheme very challenging to achieve. To this end, this article investigates for the first time the feasibility of using a compact cryogenic veto system in close vicinity to a target cryogenic detector for mitigating external particle backgrounds. The use of High Purity Germanium (HPGe) ionization detectors, which are known to have a rather low (keV) energy threshold and fast timing response, is here considered.
@@ 2 Experimental setup
The experimental setup is pictured and sketched on figure 1. It consists of two cylindrical 400 g HPGe crystals of 70 mm diameter and 20 mm height, each placed 40 mm above and below a 25 mm diameter and 25 mm height 51.7 g cylindrical crystal [7, 8]. The solid angle coverage provided by this HPGe veto arrangement is sr.
@@
The two HPGe crystals were operated in ionization mode using Aluminum (Al) electrodes evaporated on their respective top and bottom surfaces to collect the electron/hole charge carriers. These 200-nm thick electrodes were deposited by electron beam evaporation after an Argon (Ar) cleaning and passivation step using a 30-nm thick amorphous Ge:H layer [9]. The bottom crystal (BOT Ge) is equipped with planar electrodes. The top crystal (TOP Ge), which was previously used in the Edelweiss experiment [10], is equipped with ring-shaped electrodes (ID electrodes). These electrodes were specially designed to discriminate between surface and bulk events. Because this feature was unnecessary in the present work, the ring electrodes were serially connected together to mimic as closely as possible the electric field resulting from polarized planar electrodes such as those equipping the BOT Ge. The two HPGe crystals were housed within 3-mm thick copper boxes and secured using PTFE (Teflon) holders. The thermal link and the electrode signal transmission were ensured by bonding the crystals using a Kapton printed circuit board glued onto the copper box. The copper housing lids facing the LWO detector were mechanically designed to hold the LWO crystal using PTFE pieces. The HPGe detectors were polarized with a 5 V/cm electric field by biasing one electrode at +10 V and grounding the opposite one. An infra-red LED (EOLS-1650-199 from EPIGAP) was glued to the BOT Ge housing in order to regenerate the HPGe crystals [11]. The two crystals were daily illuminated for ten minutes, after which one hour was necessary to stabilize the cryostat back to base temperature. An opened and collimated 241Am source was mounted underneath the BOT Ge detector. The source was covered by a thing copper tape to stop particles.
@@
The LWO crystal was operated as a bolometer and read out with a Germanium Neutron-Transmutation-Doped (Germanium NTD) thermal sensor [12]. It was used as a target detector to test the identification and the rejection power of external particles using the HPGe ionization detectors operated in anti-coincidence mode. The NTD sensor was biased with a 1-kHz alternating constant current ( I) in order to improve the signal-to-noise ratio [13]. The full mechanical assembly was wrapped in a copper foil to shield against possible infra-red radiations. The experimental setup was operated at a temperature of 18 mK in the so-called "Actuator" cryostat at Laboratoire de Physique des Deux Infinis Irène Joliot-Curie (IJCLab, Orsay), which is a custom-made pulse-tube dilution refrigerator embedded in a single vacuum chamber (see e.g. [14] for further details). No vibration decoupling systems were used.
@@
The detectors were connected with low noise stainless steel cables. These coaxial-cables especially feature a carbon coating between the core and the shielding to dissipate the accumulation of electrostatic charges induced by mechanical vibrations, hence preserving the signal integrity and reducing microphonic noise. The electronic operation and acquisition scheme is detailed on figure 2 for both the HPGe ionization detectors and for the LWO bolometer. A first signal amplification step is performed in a cold stage using Junction Field Effect Transistors (IFN860 bi-JFETs). The output voltage signals are then filtered and further amplified through voltage amplifiers in a second electronic stage operated at room temperature. Signal acquisition was done with a commercial NI6366 X Series Data Acquisition system from National Instruments. Output signals were continuously sampled and read out at a 100 kHz frequency.
@@
Data were collected in two experimental configurations: (i) without the use (10 hours of acquisition time) and (ii) with the use (40 hours of acquisition time) of an external lead (Pb) shielding (see right panel of figure 1) placed around the cryostat. The purpose of this Pb shielding was to reduce the pile-up of events in the slow LWO detector, which features pulses of 0.5-1 s decay times and a rate of 5 Hz. Adding the Pb shielding reduced this rate by a factor 2.5. The lead shielding was made of three 5-cm thick brick walls and one partial wall of 10-cm thickness, hence partially covering the detector setup. During those two data acquisition runs, the detector setup was exposed punctually for a few hours, to a source for energy calibration and a neutron source for quenching measurements.
@@ 3 Experimental results
The operation of HPGe ionization detectors as an efficient veto system for a bolometric target detector imposes good energy and time reconstruction together with a high pulse selection efficiency down to the lowest possible threshold. This latter requirement is particularly relevant for veto detectors operated at surface, for which event rates can lead to significant pile-up probability. While a good timing resolution allows to keep accidental coincidences at acceptable rates, precise energy determination may prove to be important to better identify potential background sources present in the experimental setup environment. To meet these requirements, a fast and efficient pulse identification algorithm was developed. It takes advantage of the large difference between the detector pulse rise and decay time components, then making it almost insensitive to the presence of pile-up events. In this situation, the information about the deposited energy is mostly contained in the rise time region of the pulse. Event triggering and energy determination can then be performed in a rather small time window, hence keeping under control possible pulse distortions from pile-up effects. The present event selection and reconstruction algorithms have been applied to both the HPGe veto detectors and the LWO bolometer data streams. They are described in the following subsections.
3.1 Event selection and reconstruction
@@
The event selection and reconstruction take advantage of the HPGe and LWO detector pulse features, which exhibit a fast rise time with respect to their decay time. The amplified signals from the Ge detectors were measured to typically have an average rise time less than 30 s and a decay time typically in the 30 to 50 ms range (see figure 3). This feature is also present in the NTD sensor, which delivers much slower signals with rise times of 25 ms and decay times up to 1 s (see figure 4). As illustrated by figure 3, the bin-to-bin differentiation of such signals allows to clearly identify physical pulses above a flat baseline centered around zero. Events are thus triggered by setting an appropriate threshold above the measured baseline fluctuations. The energy associated to each detected pulse is then reconstructed by summing the amplitudes of the discretely differentiated signals over a time window covering the pulse leading edge region (i.e. non-zero derivative), and after applying an energy calibration factor. Such a procedure is equivalent to measuring the full amplitude rise of the detector signal.
For the three detectors, data were originally acquired with a 100 kHz sampling frequency. For event reconstruction, a downsampling procedure was applied to the data streams at the analysis level to reduce baseline fluctuations and to improve energy resolution. It was performed by merging and averaging the content of consecutive samples, with being the downsampling factor. The HPGe detector streams were downsampled by a factor 2, leading to a 50 kHz sampling rate. The streams of the much slower LWO detector were downsampled by a factor 250, leading to a sampling rate of 400 Hz. In both cases, this choice resulted from a trade-off between energy and time resolution. In order to reject possible spurious events in the LWO detector, pulses with less than 10-ms rise time were rejected.
The energy calibration of the three detectors was performed identifying intense gamma-ray lines originating from the decay of primordial radionuclides naturally present in the surrounding environment and from the source to which the setup has been exposed to: ( decay chain) at 510.7 keV, 583.2 keV and 2614.5 keV, annihilation at 511 keV, ( decay chain) at 609.3 keV, ( decay chain) at 911.2 keV and 969.0 keV, and at 1460.8 keV. An additional 59.5 keV gamma-ray line from the source was also used for the BOT Ge detector. The next subsections report on the characterization and performances of each detector. Results are summarized in table 1.
| Detector | Calibration | Resolution | Mean of baseline | Signal | Rate (inside |
| [V/keV] | @1461 keV [%] | r.m.s. [keV] | rise time | Pb) [Hz] | |
| TOP Ge | 651.5 0.1 | 0.6 | 1.8 | < 10 s | 8 |
| BOT Ge | 949.9 0.4 | 0.5 | 1.1 | < 10 s | 15 |
| LWO | 236.1 0.2 | 0.8 | 5.7 | 25 ms | 2 |
3.2 HPGe detector results
Figure 5(a) shows the calibration curves obtained by fitting a linear function to the identified gamma-ray peaks. Taking into account the gain of 1000 provided by the amplification stage, the BOT Ge detector featuring planar electrodes is found to have a nV/keVee sensitivity while a nV/keVee sensitivity is measured for the TOP Ge detector having the ring-shaped electrodes. Non-linearities in the energy scale below 3 MeVee are found to be negligible.
Relative energy resolutions measured at the previously mentioned gamma-ray line energies are displayed on figure 5(b). They are found to be better than 1 % above 400 keV and are very similar for both detectors. They are well described by the characteristic energy dependence as expected from Ge semiconductor detectors [15]:
@@ 3.2 HPGe detector results
Electronic noise contribution for each of the HPGe detector was independently characterized by sampling the baseline fluctuations during each data run. The mean of the distribution of the baseline r.m.s. values were measured to be 1.1 keV and 1.8 keV for the BOT and TOP Ge detectors, respectively. These measurements come in very good agreement with the previously mentioned fit results (see figure 5 (b)). The larger value, in keVee units, observed for the TOP Ge detector is mostly due to its smaller sensitivity. Some spurious periodic pulses were also identified and always reconstructed with energies below 15 keV. After investigation, those were attributed to cross-talk effects originating from the 1-kHz modulation current flowing through the LWO NTD sensor. Since the sampling and modulation clocks were synchronous, this cross-talk noise could be mitigated by removing very low energy pulses that were identified in-time with the leading edges of the NTD current modulation signal. This requirement introduced a small inefficiency of at most a few percents in the selection of events below 15 keV.
With the noise treatment described above, detection thresholds of 5 keV and 10 keV were applied to select and to reconstruct events in the BOT and TOP Ge detectors, respectively. The corresponding counting rates, integrated over the full energy spectra up to 20 MeV, were measured to be about 15 and 8 counts/s, respectively, during a typical run using the Pb shielding. The difference in rate between the two detectors can be mostly accounted for by the activity of the source. Figure 6 further illustrates this difference, where the energy spectra of the selected and reconstructed events in the TOP and BOT Ge detectors are compared over the full 20 MeV range and in the region below 600 keV. With no Pb shielding surrounding the cryostat, the rates were found to be significantly higher, ranging between 37-43 and 26-31 counts/s for the BOT and TOP Ge detectors, respectively.
@@
As mentioned previously, good time resolution of the Ge veto detectors is necessary in order to limit accidental rejection of good candidates in the LWO bolometer. The arrival time of an event in the BOT and TOP Ge detectors was determined to better than 10 s, owing to the short rise time of pulses. Given the counting rates, true coincidences between the two Ge detectors could clearly be observed with a negligible level of accidentals. This is illustrated also in Figure 6, where the TOP and BOT Ge energy spectra are also shown for events detected in coincidence (orange and green curves). With the exception of the 511 keV annihilation line, all other X-ray and gamma-ray lines corresponding to full absorption peaks disappear. Moreover, a large fraction of events detected in coincidence is located in the high energy portion of the spectra. These events come from atmospheric muons crossing both detectors. The rate of coincidences between the TOP and BOT Ge detectors for a typical background run using the Pb shielding is 0.8 counts/s.
3.3 LWO bolometer results
The calibration curve obtained by fitting a linear function to the identified gamma-ray peaks in the LWO detector is shown in Figure 5(a). Taking into account a gain of 8000 from the NTD signal warm amplification stage, the sensitivity of the LWO bolometer was determined to be nV/keV. Non-linearities in the energy scale below 3 MeV were found to be negligible. Figure 8 shows the energy spectrum measured with the LWO bolometer during a run performed with the Pb shielding. In addition to environmental gamma-ray lines, a peak at MeV due to the neutron capture reaction is visible. The peak intensity is here enhanced by the use of a neutron source. It is shifted by with respect to the 4.78 MeV total energy released in the reaction because of the so-called thermal quenching effect. This effect is caused by the reduced scintillation yield of alpha and tritron particles with respect to electrons or gamma rays, which consequently lead to larger heat signals. This effect is expected to be of the order of a few percents in Lithium-containing scintillating bolometers [7].
The measured energy resolution below 6 MeV is shown on figure 5 (b). It is well-described by the following relation [16, 17]: